If twice the square of a positive number is added to the product of -1 and the number,the result is 45.Find the number.
Answer by faceoff57(108)

2x^2+-1x=45
2x^2-x-45=0
(2x+9)(x-5)=0
2x+9=0
2x=-9
x=-9/2
OR x-5=0 x=5
THE NUMBER IS 5